Supplemental Operations Report 25-DEC-2022

Supplemental Operations Report 25-DEC-2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Power system programing

Action taken for non-nominal systems: Nothing yet. I will go through and reprogram it in the next few days.

Generator: Running from about 8 pm to 8 am. The generator was repaired on Thursday. The problem was the governor, which had lost a small plastic cotter pin and as a result was unstable and the vibration of the engine was causing the surging. How crazy! We also needed the oil pressure changed as it was faulty and we had the bushings replaced because they would have needed it soon. The technician was great and spent a lot of time showing us what things were that would commonly need replaced and how to replace them. I am very grateful to not have to worry every night about losing power!

ScienceDome Dual Split: on at 65 degrees overnight (when I remember). I need to find the manual in order to program it but I don’t know where it is.

Solar— Nominal but router needs reprogrammed

Solar— SOC Last 24 hours:



Avg not showing the way it is currently operating

Propane Reading, station tank – 61 %

Propane Reading, director tank— 75 %

Propane Reading, intern tank— 72 %

Propane Reading, generator— 60 %

Ethanol Free Gasoline – 0 gallons

Water (static tank) – not checked gallons

Water in GreenHab – ~150 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 350 gallons

Hab toilet tank emptied: See crew operations reports

Sojourner rover used: no

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town

CrewCar used and why, where? Yes, to town

Luna used and why, where? Yes, to town

General notes and comments: We have Christmas lights up outside for the first time at the station. They are a simple string across three rovers outside the RAM but they are quite festive and cheerful.

Summary of internet: Nominal.

Summary of suits and radios: See crew operations reports.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: I was unable to completely unclog the toilet but over several fills and flushes with hot water and lots of dish soap, I made enough progress for us to get through the next two crews and then January, during the time we have no crew, we will replace the toilet tank with a new one and plumb in a cleanout value.

Summary of Outpost operations: The deck with the clothesline has been cleaned off in anticipation of required repairs. We need to replace the surface material and reinstall the clothesline. General outside cleanup.

Summary of GreenHab operations: Supplemental light 10-2 pm. The beans produced and were so good we planted more beans. Beans, peas and tomatoes are all flowering.

Summary of ScienceDome operations: We repaired the leak near the north window.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Peter and Crew Astronomer worked on the Robotic Observatory last Sunday.

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Merry Christmas!

Copyright © The Mars Society. All rights reserved. | Main Site